A company wanted to determine what percentage of its employees stayed for at least 1, 2, and 3 years. The data they compiled is given below and covers 200 male employees and 300 female employees.
1 year 2 years 3 years
Male 0.67 0.45 0.20
Female 0.73 0.64 0.39


What is the probability that an employee stayed for at least one year? Round your answer to the nearest hundredth.

Answer by Theo(13342)

percents are:

	     1 year     2 years     3 years

Male           0.67        0.45        0.20
Female         0.73        0.64        0.39



there are 200 males and there are 300 females.

.67 * 200 = 134 males who stayed at least one year.
.73 * 300 = 219 females who stayed at least one year.

total is 353 out of 500 employees who stayed as least 1 year.

353 / 500 = .706

the probability that an employee stayed at least one year is therefore equal to .706.